OpenStudy (anonymous):

Separable equations \[\frac{dy}{dx}=\frac{ycosx}{1+y^{2}}\] \[\int \frac{1+y^{2}}{y}dy= \int cosx dx\] \[\int \frac{1+y^{2}}{y}dy= sinx +C\]

OpenStudy (anonymous):

u=1+y^2 ?

OpenStudy (anonymous):

simplify the integrand on the left side by doing long division first....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y+ 1/y

OpenStudy (anonymous):

\[\huge \frac{1+y^2}{y}=\frac{1}{y}+\frac{y^2}{y}=y^{-1}+y \]

OpenStudy (anonymous):

yep... that's easily integrated huh?

OpenStudy (anonymous):

yep y^2 /2

OpenStudy (anonymous):

what about y^-1 ?

OpenStudy (anonymous):

1+ y^2 /2

OpenStudy (anonymous):

1+ (y^2)/2

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

\[\large \int \frac{1}{y}dy=lny+C \]

OpenStudy (anonymous):

duh...thanks

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

yw...:)